OpenStudy (anonymous):

Can someone check my radical answers.... 8√7 / 7√8= 4/7??? (-8-4√5)(2-√5)=4???

OpenStudy (anonymous):

@satellite73 <--- lol u always help me

OpenStudy (johnweldon1993):

Hmm...can you show your work for #1?

OpenStudy (jdoe0001):

and #2 :)

OpenStudy (jdoe0001):

\(\bf \cfrac{8\sqrt{7}}{7\sqrt{8}}\ne \cfrac{4}{7}\)

OpenStudy (anonymous):

(-8-4√5) (2-√5)= -16+8√5-8√5+4√25=-16+8√5-8√5+20= 4

OpenStudy (anonymous):

thats for #2

OpenStudy (johnweldon1993):

Yeah that looks correct to me

OpenStudy (johnweldon1993):

but what about #1?

OpenStudy (anonymous):

jdoe answered it and that was exactly how i did so thanks!

OpenStudy (johnweldon1993):

Oh he already answered it? That's odd..I'm sorry jdoe but I'd have to disagree with your answer...for now

OpenStudy (anonymous):

ok then i am confused

OpenStudy (anonymous):

is it √14/7